Circular No. 7246

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

CASSIOPEIA A H. Tananbaum, Harvard-Smithsonian Center for Astrophysics (CfA),

on behalf of the Chandra X-ray Observatory (representing scientists at the Chandra X-ray Center, Goddard Space Flight Center, CfA, Marshall Space Flight Center, Massachusetts Institute of Technology, Pennsylvania State University, and Space Research Organization of the Netherlands), reports: “The team has determined the following position for the pointlike source seen near the center of Cas A during the Chandra First Light observation: α = 23^{h}23^{m}27^{s}.94, δ = +58^{o}48′42′′.4 (ICRS, equinox J2000.0). The uncertainty in the x-ray position is currently ∼ 2′′.5 in radius, but is expected to be reduced to ∼ 1′′ as the calibration proceeds. A quick look at the x-ray data indicates that any extent is < 2′′. The count rate for this source (with a backside-illuminated ACIS chip) is ∼ 1/3000 of the total x-ray emission from Cas A over the 0.2–10-keV Chandra band. An accurate conversion to flux and luminosity requires the in-flight spectral response matrices and a careful spectral fitting, but the x-ray photometry alone suggests that the source luminosity is at least 10^{32} erg/s, if it is at the same distance as Cas A. The source lies within 5′′ of the ‘expansion center’ for the remnant determined by van den Bergh and Pritchet (1986, Ap.J. 307, 723) after precessing their position to equinox J2000.0. There are no obvious pointlike counterparts within 5′′ on the Very Large Array 20-cm

maps (http://www.nrao.edu/vla/html/VLA-images.shtml), in the van

den Bergh-Pritchet R-band image (limiting mag ∼ 23.5), or in the [S II] image of Fesen and Gunderson (1996, Ap.J. 470, 967). Longer exposures with Chandra are planned, and these should allow a search for periodic pulsations. We suggest that this object might well be the collapsed star formed at the time of the Cas A supernova explosion.”

NOVA CIRCINI 1999 Visual magnitude estimates: Aug. 25.459 UT, 8.1 (A. Pearce, Ned-

lands, W. Australia); 25.617, 8.2 (Pearce); 25.996, 8.9 (J. G. de S. Aguiar, Campinas, Brazil); 26.938, 8.8 (H. C. Vital, Rio de Janeiro, Brazil); 27.490, 9.2 (Pearce); 28.897, 9.1 (Vital); 29.013, 9.1 (Vital); 30.529, 9.8 (Pearce); 30.901, 9.5 (Vital); 31.533, 9.9 (Pearce); 31.933, 9.0 (Vital); Sept. 1.088, 9.4 (Vital); 1.897, 10.0 (Vital); 2.037, 10.1 (Vital).

1999 September 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green